I need help in getting the answer.
One side of a rectangular stage is 2 meters longer than the other. If the diagonal is 10 meters,then what are the lengths of the sides?

Answer by checkley77(12844)

L=W+2
L^2+W^2=10^2
(W+2)^2+W^2=100
W^2+4W+4+W^2=100
2W^2+4W+4-100=0
2W^2+4W-96=0
2(W^2+2W-48)=0
2(W-6)(W+8)=0
W-6=0
W=6 METERS ANS. FOR THE WIDTH.
L=6+2=8 METERS ANS. FOR THE LENGTH,
PROOF:
6^2+8^2=100
36+64=100
100=100